Solve the following partial differential equations 


A) (y-z)∂z/∂x +(z-x)∂z/∂y=(x-y)


B) ∂^2z/∂x^2 +7 ∂^2z/∂x∂y + 6 ∂^2z/∂y^2 =sin(x+2y)


1
Expert's answer
2021-02-24T06:07:54-0500

(A) Given equation is-


"(y\u2212z)p+(z\u2212x)q=x\u2212y"


This is Lagrange's linear equation "Pp+Qq=R"


∴ The auxiliary equation is

   

    "\dfrac{dx}{y-z}=\dfrac{dy}{z-x}=\dfrac{dz}{x-y}"


Each is equal to "\dfrac{dx+dy+dz}{y+z+z-x+x-y}"


"=\dfrac{d(x+y+z)}{0}"




d(x+y+z)=0


∴ On Integration, x+y+z=a


Also, each ratio is equal to


"\Rightarrow d(x+y+z)=0"


"\dfrac{xdx+ydy+zdz}{x(y−z)+y(z−y)+z(x−y)}"


"=\dfrac{\dfrac{1}{2}d(x^2+y^2+z^2)}{0}"


"\Rightarrow d(x^2+y^2+z^2)=0"


Integrating, "x^2+y^2+z^2=b"


∴ The general solution is


"\phi(x+y+z,x^2+y^2+z^2)=0"


(B )We have given,


"\dfrac{d^2z}{dx^2}+\dfrac{7d^2z}{dxdy}+\dfrac{6d^2z}{dy^2}=sin(x+2y)"


Then the auxiliary equation will be

"(m^2+7m+6)=0\\

m^2+m+6m+6=0\\

m(m+1)+6(m+1)=0\\

m=-1,-6"


Hence, "C.F=f1(y-x)+f2(y-6x)"


Now, Particular integral can be calculated as


"P.I= \dfrac{sin(x+2y)}{(D^2+7DD'+D'^2)}"


"= \dfrac{sin(x+2y)}{(-1)+7(-2)+6(-24)}"


"=-\dfrac{sin(x+2y)}{39}"


Therefore, the solution of partial differential equation is "C.F+P.I"


"=f1(y-x)+f2(y-6x)-\dfrac{sin(x+2y)}{39}"